### v1

@@ Appendix E Synthetic data examples: hyperparameters and sample data sets
Here, we provide specific details on the data generating mechanisms S1-S4 used in Section D.2, including hyperparameter configurations and sample draws from each parent distribution. The input space is $\mathcal{X}=[0,10]$, i.e., $d=1$ and elements of the input space are scalars.
@@
Here, recall that $C_{y}(x,x^{\prime})=\sigma^{2}\mathcal{M}_{\nu}(|x-x^{\prime}|)$. We specify $\sigma^{2}=1$, $\nu=2.5$, and the Matérn length-scale parameter is $\rho=0.5$. The training data include additive Gaussian white noise with variance $\tau^{2}=0.1$. Sample draws from a mean-zero Gaussian process with corresponding kernel are shown in Figure E.1.
@@
Here, $C_{y}(x,x^{\prime})=\sigma^{2}f_{0}(x,x^{\prime};r_{0})$, where $f_{0}$ is as in Eq. 2.1. We specify $\sigma^{2}=1$ and $r_{0}=1.5$. The training data include additive Gaussian white noise with variance $\tau^{2}=0.1$. Sample draws from a mean-zero Gaussian process with corresponding kernel are shown in Figure E.1.
@@
The training data include additive Gaussian white noise with variance $\tau^{2}=0.475$. Sample draws from a mean-zero Gaussian process with corresponding kernel are shown in Figure E.1.
@@
The training data include additive Gaussian white noise with variance $\tau^{2}=0.506$. Sample draws from a mean-zero Gaussian process with corresponding kernel are shown in Figure E.1.
Figure E.1: Four sample draws from each of the four parent distributions described in Table D.1.
Appendix F Prior mean and kernel for temperature data
F.1 Spatial modeling of data time slices
Prior mean function for time slices. For time-slices of daily maximum temperature measurements, there are two primary features to account for in the prior mean: relationships with orography and distance from the coast. Specifically, we generally expect (1) inverse relationships between temperature and elevation (via so-called “lapse-rate” theory); and (2) direct relationships between temperature and distance to coastlines. We can furthermore include additional topographic variables of slope and aspect, which respectively describe the direction in which each location faces (ranging from zero to $2\pi$) and the steepness. Collectively, slope and aspect account for the effect of solar radiation and exposure on daily maximum temperatures. The elevation is obtained from the 800m digital elevation map used to generate the PRISM data product; slope and aspect are calculated using the raster package for R [61]. The distance-to-coast (km) variable is calculated by NASA’s Ocean Biology Processing Group [62] at a global grid of $1/16^{\circ}$ using the Generic Mapping Tools package [63] (note that this variable only calculates distances to ocean coastlines and not, e.g, other landlocked bodies of water). Together, the prior mean function is a linear function of these covariates:
$\mu({\bf s})=\alpha_{0}+\alpha_{1}\text{Elev}({\bf s})+\alpha_{2}\text{Slope}({\bf s})+\alpha_{3}\text{Aspect}({\bf s})+\alpha_{4}\text{Dist2Coast}({\bf s}),$
where we update the coefficients $\{\alpha_{j}:j=0,\dots,4\}$ within the MCMC.
@@ F.1 Spatial modeling of data time slices
$C_{\text{core}}({\bf s},{\bf s}^{\prime};{\boldsymbol{\theta}}_{\text{core}})=\sigma({\bf s})\sigma({\bf s}^{\prime})\frac{\left|{\bf\Sigma}({\bf s})\right|^{1/4}\left|{\bf\Sigma}({\bf s}^{\prime})\right|^{1/4}}{\left|\frac{{\bf\Sigma}({\bf s}^{\prime})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right|^{1/2}}\mathcal{M}_{1.5}\left(\sqrt{({\bf s}-{\bf s}^{\prime})^{\top}\left(\frac{{\bf\Sigma}({\bf s})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right)^{-1}({\bf s}-{\bf s}^{\prime})}\right),$ (F.1)
where $\mathcal{M}_{1.5}$ is the Matérn kernel with smoothness equal to 1.5. We parameterize the spatially-varying signal variance as
@@
${\bf\Sigma}({\bf s})={\bf\Gamma}({\bf s}){\bf\Lambda}({\bf s}){\bf\Gamma}({\bf s})^{\top},\hskip 12.91663pt\text{where}\hskip 12.91663pt{\bf\Lambda}({\bf s})=\left[\begin{array}[]{cc}\lambda_{1}({\bf s})&0\\ 0&\lambda_{2}({\bf s})\end{array}\right],{\bf\Gamma}({\bf s})=\left[\begin{array}[]{cc}\cos\gamma({\bf s})&-\sin\gamma({\bf s})\\ \sin\gamma({\bf s})&\cos\gamma({\bf s})\end{array}\right],$
@@
$\begin{array}[]{c}\log\lambda_{1}({\bf s})=\omega^{\lambda_{1}}_{0}+\omega^{\lambda_{1}}_{1}\text{Elev}({\bf s})+\omega^{\lambda_{1}}_{2}\text{Dist2Coast}({\bf s}),\\[2.15277pt] \log\lambda_{2}({\bf s})=\omega^{\lambda_{2}}_{0}+\omega^{\lambda_{2}}_{1}\text{Elev}({\bf s})+\omega^{\lambda_{2}}_{2}\text{Dist2Coast}({\bf s}),\\[2.15277pt] \log\frac{\frac{2}{\pi}\gamma({\bf s})}{1-\frac{2}{\pi}\gamma({\bf s})}=\omega^{\gamma}_{0}+\omega^{\gamma}_{1}\text{Elev}({\bf s})+\omega^{\gamma}_{2}\text{Dist2Coast}({\bf s}).\end{array}$ (F.2)
@@
These specific transformations are chosen so that the coefficients $\{\omega_{j}^{(\cdot)}\}$ can be unrestricted and still yield valid ${\bf\Sigma}(\cdot)$. For the sparsity-inducing kernel, the primary choice is to specify $n_{1}$ and $n_{2}$, the number of sets ($n_{1}$) of correlated yet distance-unrelated regions ($n_{2}$). Following the guidance of Section D.1, we err on the side of choosing these to be larger than we might expect. We consider three configurations, each with 100 total bump functions: (1) $n_{1}=2$ and $n_{2}=50$; (2) $n_{1}=4$ and $n_{2}=25$; and (3) $n_{1}=10$ and $n_{2}=10$. Note that in this implementation, for the sparse kernel, we use an anisotropic version of the Wendland kernel that has separate $x$- and $y$-coordinate truncation radii. Furthermore, we consider three upper limits on the Wendland truncation radii (5, 10, and 15 units of degrees longitude/latitude) since we found that otherwise the truncation radii tend to increase without bound in the training. Limiting these radii to be relatively small ensures that the resulting kernel has the chance to impose sparsity on the data.
@@
In summary, we have five hyperparameters for the prior mean function ($\{\alpha_{j}:j=0,\dots,4\}$), twelve hyperparameters for the core kernel (${\boldsymbol{\theta}}_{\text{core}}=\{\phi_{j},\omega_{j}^{\lambda_{1}},\omega_{j}^{\lambda_{2}},\omega_{j}^{\gamma}:j=0,1,2\}$), one hyperparameter for the noise variance. In the sparse kernel, 100 bump functions means that we have $5033+n_{1}n_{2}(d+3)=3+100\times 5=$ 503 hyperparameters. Overall, this yields $5+12+1+503=521$ hyperparameters for the GP applied to each time slice.
F.2 Spatio-temporal modeling
@@
Prior mean function in space and time. In light of the fact that we are modeling high-dimensional space-time data and our kernel can discover and capitalize upon sparsity, it is to our benefit to propose a prior mean function that explains as much of the known spatial and temporal structure in the data as possible. The most prominent structure in a data set involving temperature is the annual cycle, wherein daily maximum temperatures in the Northern Hemisphere are generally lowest in the winter and highest in the summer. As in Section 4.1, there is an underlying geospatial climatology to measurements of temperature including an implicit zonal structure wherein spatial locations farther from the equator are generally cooler, a strong influence from orographic variability, and the influence of coastlines. To account for each of these features, we propose the following prior mean function for each ${\bf x}=({\bf s},t)\in\mathcal{X}=\mathcal{S}\times\mathcal{T}$:
$\mu({\bf x})=\alpha_{0}({\bf s})+\sum_{j=1}^{J}\alpha_{j}({\bf s})v_{j}(t).$ (F.3)
@@
$\alpha_{j}({\bf s})=\sum_{m=1}^{M_{s}}\delta_{jm}^{s}q^{s}_{m}({\bf s})+\sum_{m=1}^{M_{e}}\delta_{jm}^{e}q^{e}_{m}(e)+\sum_{m=1}^{M_{c}}\delta_{jm}^{d}q^{c}_{m}(c).$ (F.4)
@@
The linearity assumptions of Equations F.3 and F.4 allow us to write the mean function $\mu({\bf x})$ as a linear function of a set of covariates, as denoted in Section 3.1. Let ${\bf W}_{\text{time}}$ denote a $N_{\text{time}}\times(J+1)$ matrix of cubic spline functions ($J$ of which correspond to the cubic splines and 1 for an intercept), where $N_{\text{time}}=|\mathcal{T}|$, and let ${\bf W}_{\text{space}}$ denote a $N_{\text{space}}\times(1+M_{s}+M_{e}+M_{c})$ matrix of thin plate spline functions with an intercept, where $N_{\text{space}}=|\mathcal{S}|$. The design matrix ${\bf W}$ can then be written as
${\bf W}={\bf W}_{\text{space}}\otimes{\bf W}_{\text{time}}$ (F.5)
@@
(Here we suppose $N=N_{\text{space}}N_{\text{time}}$, although any inputs ${\bf x}$ with missing temperature measurements are omitted.) The matrix ${\bf W}$ in Equation F.5 is then $N\times[(J+1)\times(1+M_{s}+M_{e}+M_{c})]$, meaning that this yields $(J+1)(1+M_{s}+M_{e}+M_{c})$ hyperparameters for the prior mean. We re-label the coefficients $\{\delta^{(\cdot)}_{jm}\}$ from Equation F.4 into an appropriately ordered vector ${\boldsymbol{\beta}}$ such that the prior mean function across all $N$ measurements can be concisely written as ${\bf W}{\boldsymbol{\beta}}$, as used in Section 3.1.
@@
Figure F.1: Comparison of the fitted prior mean (using ordinary least squares) versus input data for four selected stations (shown in panel a.) across all time points (panel b.), where the raw data are shown in black and the fitted prior mean in red. Panel c. then shows spatial maps of the 2-, 5-, and 10-day lagged autocorrelation from the input data (top) and residuals (data minus OLS prior mean; bottom), while panel d. shows the length-scale and signal variance from semivariograms fitted separately for each for the input data (top) and residuals (data minus OLS prior mean; bottom).
@@
After some trial and error and in order to maintain a compromise between sufficient flexibility and overfitting, we fix $J=8$, $M_{s}=200$, $M_{e}=5$, and $M_{c}=5$, resulting in $1899$ hyperparameters for the prior mean. To visualize the selected prior mean function, we show the fitted prior mean – for starters using ordinary least squares (OLS), but later these hypereparameters will be trained online with the GP – for specific stations and days versus the corresponding raw temperature measurements; see Figure F.1. Across all input locations, the OLS linear mean explains 94.1% of the variability in the daily temperature data, which is impressive since there is one hyperparameter for every $\approx 1100$ data points. For the specific stations shown in Figure F.1(b), it is clear that the seasonal cycle appropriately captures the general trend of the data, such that the spatially-varying nature of the coeffients can describe different amplitudes and even shapes in the the seasonal cycle: note that the daily maximum temperatures range from approximately 5C to 35C at the Matador station and approximately -10C to 30C at the Sibley station; furthermore, the seasonal cycle is extremely smooth at the Minden station but “bumpy” at Morehead City.
@@
It is important to note how the variability explained by a flexible prior mean function “increases” the sparsity that can be leveraged by the kernel. In the time dimension, Figure F.1(c) shows how the 2-, 5-, and 10-day lagged autocorrelations at each gauged location differ for the input data versus the residuals (the input data minus the OLS prior mean function). Even at a lag of 10 days, the data show strong autocorrelations, in excess of 0.6. However, when removing the fitted OLS prior mean, 10-day lagged autocorrelations essentially vanish; even for 5-day lags, the autocorrelations are quite small. Spatially, Figure F.1(d) shows the length-scale and signal variance from a variogram fitted to either the input data (black) or the OLS residuals separately for each day. Here, the effect of the prior mean function is less clear: in particular, the length-scale of the residuals appear to be on a similar order of magnitude as the input data. On the other hand, the variability of the residuals is significantly reduced relative to the input data. In summary, it is clear that the prior mean function allows the data to be quite sparse in the time domain (i.e., correlations beyond ±$\pm 5$ days can be safely ignored) and perhaps less sparse in the spatial domain. These differences in the structure explained by the prior mean (temporal vs. spatial) have important implications for how we set up the kernel.
@@
Kernel. For the full space-time data, we use a core kernel similar to the one in Equation F.1, albeit generalized to a three-dimensional input space, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})$. The anisotropy matrix again uses the eigendecomposition ${\bf\Sigma}({\bf x})={\bf\Gamma}({\bf x}){\bf\Lambda}({\bf x}){\bf\Gamma}({\bf x})^{\top}$, where
${\bf\Lambda}({\bf x})=\left[\begin{array}[]{ccc}\lambda_{1}({\bf s})&0&0\\ 0&\lambda_{2}({\bf s})&0\\ 0&0&\lambda_{3}(t)\end{array}\right],\hskip 12.91663pt{\bf\Gamma}({\bf x})=\left[\begin{array}[]{ccc}\cos\gamma({\bf s})&-\sin\gamma({\bf s})&0\\ \sin\gamma({\bf s})&\cos\gamma({\bf s})&0\\ 0&0&1\end{array}\right].$
@@
The formulas for $\lambda_{1}({\bf s})$, $\lambda_{2}({\bf s})$, and $\gamma({\bf s})$ are as in Equation F.2, and the length-scale in time is a constant, i.e., $\lambda_{3}(t)\equiv\lambda_{3}$. Note that in this case, the term inside the Matérn correlation is
@@
For the sparse kernel, the Wendland component is again anisotropic with coordinate-specific length scales (and coordinate-specific truncation limits); following Section 4.1 we use 10 units for the spatial truncation limits and 5 days for the temporal truncation limits based on the fact that the prior mean function explains autocorrelation beyond 5 days for a large majority of the domain. We again use 100 bump functions with $n_{1}=4$ and $n_{2}=25$; however, again based on the fact that the prior mean explains so much of the temporal variability, we restrict the bump functions to live in the spatial domain only. In total, this yields $521+1+1=523$ hyperparameters for the kernel (the same 521 from the time-slice modeling in Section 4.1 plus $\lambda_{3}$ and the Wendland length-scale in the time dimension).
Appendix G Supplemental figures
Figure G.1: Sample bump functions $g({\bf x})$ from Equation 1 for different values of the shape parameter $b$ with fixed amplitude $a$ and radius $r$. As $b\rightarrow 0$, the bump function becomes uniform on $(-r,r)\subset\mathbb{R}^{1}$; as $b\rightarrow\infty$, the bump functions become delta function centered at $\mathbf{h}$. Note that the bump functions are exactly zero for distances from the centroid ${\bf h}$ greater than or equal to $r$.
Figure G.2: Posterior summaries of hyperparameters for the January 14, 2001 time slice of daily maximum temperature. Panels (a), (b), and (c) show the prior mean function, square root of the signal variance, and anisotropy length-scales, calculated using the posterior median hyperparameters and the physical covariates. Panel (d) shows implied correlations between eight reference locations and all other points in the input domain.

### v3

@@ Appendix D Synthetic data examples: hyperparameters and sample data sets
Here, we provide specific details on the data generating mechanisms S1-S4 used in Section 4.2, including hyperparameter configurations and sample draws from each parent distribution. The input space is $\mathcal{X}=[0,10]$, i.e., $d=1$ and elements of the input space are scalars. All cases include additive Gaussian white noise wth variance equal to 20%, 10%, and 5% of the signal variance.
@@
Here, recall that $C_{y}(x,x^{\prime})=\sigma^{2}\mathcal{M}_{\nu}(|x-x^{\prime}|)$. We specify $\sigma^{2}=1$, $\nu=2.5$, and the Matérn length-scale parameter is $\rho=0.5$. Sample draws from a mean-zero Gaussian process with corresponding kernel and 10% noise variance are shown in Figure D.1.
@@
Here, $C_{y}(x,x^{\prime})=\sigma^{2}f_{0}(x,x^{\prime};r_{0})$, where $f_{0}$ is as in Eq. 2.1. We specify $\sigma^{2}=1$ and $r_{0}=1.5$. Sample draws from a mean-zero Gaussian process with corresponding kernel and 10% noise variance are shown in Figure D.1.
@@
Sample draws from a mean-zero Gaussian process with corresponding kernel and 10% noise variance are shown in Figure D.1.
@@
Sample draws from a mean-zero Gaussian process with corresponding kernel and 10% noise variance are shown in Figure D.1.
Figure D.1: Four sample draws from each of the four parent distributions described in Table 1.
Appendix E Prior mean and kernel for daily maximum temperature data
E.1 Spatial modeling of daily data
Prior mean function for daily data . In order to focus on the second-order properties of daily maximum temperature, for the spatial-only modeling we elect to use a spatially-constant prior mean function. We revisit this decision by proposing a more flexible prior mean function for space-time modeling in Section E.2.
@@ E.1 Spatial modeling of daily data
$C_{\text{core}}({\bf s},{\bf s}^{\prime};{\boldsymbol{\theta}}_{\text{core}})=\sigma({\bf s})\sigma({\bf s}^{\prime})\frac{\left|{\bf\Sigma}({\bf s})\right|^{1/4}\left|{\bf\Sigma}({\bf s}^{\prime})\right|^{1/4}}{\left|\frac{{\bf\Sigma}({\bf s}^{\prime})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right|^{1/2}}\mathcal{M}_{1.5}\left(\sqrt{({\bf s}-{\bf s}^{\prime})^{\top}\left(\frac{{\bf\Sigma}({\bf s})+{\bf\Sigma}({\bf s}^{\prime})}{2}\right)^{-1}({\bf s}-{\bf s}^{\prime})}\right),$ (E.1)
where $\mathcal{M}_{0.5}$ is the Matérn kernel with smoothness equal to 1.5. We parameterize the spatially-varying signal variance as
@@
${\bf\Sigma}({\bf s})={\bf\Gamma}({\bf s}){\bf\Lambda}({\bf s}){\bf\Gamma}({\bf s})^{\top},\qquad\text{where}\qquad{\bf\Lambda}({\bf s})=\left[\begin{array}[]{cc}\lambda_{1}({\bf s})&0\\ 0&\lambda_{2}({\bf s})\end{array}\right],{\bf\Gamma}({\bf s})=\left[\begin{array}[]{cc}\cos\gamma({\bf s})&-\sin\gamma({\bf s})\\ \sin\gamma({\bf s})&\cos\gamma({\bf s})\end{array}\right],$
@@
$\begin{array}[]{c}\log\lambda_{1}({\bf s})=\omega^{\lambda_{1}}_{0}+\omega^{\lambda_{1}}_{1}\text{Elev}({\bf s})+\omega^{\lambda_{1}}_{2}\text{Dist2Coast}({\bf s}),\\[2.15277pt] \log\lambda_{2}({\bf s})=\omega^{\lambda_{2}}_{0}+\omega^{\lambda_{2}}_{1}\text{Elev}({\bf s})+\omega^{\lambda_{2}}_{2}\text{Dist2Coast}({\bf s}),\\[2.15277pt] \log\frac{\frac{2}{\pi}\gamma({\bf s})}{1-\frac{2}{\pi}\gamma({\bf s})}=\omega^{\gamma}_{0}+\omega^{\gamma}_{1}\text{Elev}({\bf s})+\omega^{\gamma}_{2}\text{Dist2Coast}({\bf s}).\end{array}$ (E.2)
@@
These specific transformations are chosen so that the coefficients $\{\omega_{j}^{(\cdot)}\}$ can be unrestricted and still yield valid ${\bf\Sigma}(\cdot)$. For the sparsity-inducing kernel, the primary choice is to specify $n_{1}$ and $n_{2}$, the number of sets ($n_{1}$) of correlated yet distance-unrelated regions ($n_{2}$). Following the guidance of Section 4.1, we err on the side of choosing these to be larger than we might expect. We consider three configurations, each with 100 total bump functions: (1) $n_{1}=2$ and $n_{2}=50$; (2) $n_{1}=4$ and $n_{2}=25$; and (3) $n_{1}=10$ and $n_{2}=10$. Note that in this implementation, for the sparse kernel, we use an anisotropic version of the Wendland kernel that has separate $x$- and $y$-coordinate truncation radii. Furthermore, we consider three upper limits on the Wendland truncation radii (5, 10, and 15 units of degrees longitude/latitude) since we found that otherwise the truncation radii tend to increase without bound in the training. Limiting these radii to be relatively small ensures that the resulting kernel has the chance to impose sparsity on the data.
@@
In summary, we have five hyperparameters for the prior mean function ($\{\alpha_{j}:j=0,\dots,4\}$), twelve hyperparameters for the core kernel (${\boldsymbol{\theta}}_{\text{core}}=\{\phi_{j},\omega_{j}^{\lambda_{1}},\omega_{j}^{\lambda_{2}},\omega_{j}^{\gamma}:j=0,1,2\}$), one hyperparameter for the noise variance. In the sparse kernel, 100 bump functions means that we have $3+n_{1}n_{2}(d+3)=3+100\times 5=503$ hyperparameters. Overall, this yields $5+12+1+503=521$ hyperparameters for the GP applied to each time slice.
E.2 Spatio-temporal modeling
@@
Prior mean function in space and time. In light of the fact that we are modeling high-dimensional space-time data and our kernel can discover and capitalize upon sparsity, it is to our benefit to propose a prior mean function that explains as much of the known spatial and temporal structure in the data as possible. The most prominent structure in a data set involving temperature is the annual cycle, wherein daily maximum temperatures in the Northern Hemisphere are generally lowest in the winter and highest in the summer. Also, there is an underlying geospatial climatology to measurements of daily maximum temperature including an implicit zonal structure wherein spatial locations farther from the equator are generally cooler, a strong influence from orographic variability, and the influence of coastlines. Specifically, we expect (1) inverse relationships between temperature and elevation (via so-called “lapse-rate” theory); and (2) direct relationships between temperature and distance to coastlines. Elevation is obtained from the 800m digital elevation map used to generate the PRISM data product; slope and aspect are calculated using the raster package for R [65]. The distance-to-coast (km) variable is calculated by NASA’s Ocean Biology Processing Group [66] at a global grid of $1/16^{\circ}$ using the Generic Mapping Tools package [67] (note that this variable only calculates distances to ocean coastlines and not, e.g, other landlocked bodies of water). To account for each of these features, we propose the following prior mean function for each ${\bf x}=({\bf s},t)\in\mathcal{X}=\mathcal{S}\times\mathcal{T}$:
$\mu({\bf x})=\alpha_{0}({\bf s})+\sum_{j=1}^{J}\alpha_{j}({\bf s})v_{j}(t).$ (E.3)
@@
$\alpha_{j}({\bf s})=\sum_{m=1}^{M_{s}}\delta_{jm}^{s}q^{s}_{m}({\bf s})+\sum_{m=1}^{M_{e}}\delta_{jm}^{e}q^{e}_{m}(e)+\sum_{m=1}^{M_{c}}\delta_{jm}^{d}q^{c}_{m}(c).$ (E.4)
@@
The linearity assumptions of Equations E.3 and E.4 allow us to write the mean function $\mu({\bf x})$ as a linear function of a set of covariates, as denoted in Section 3.1. Let ${\bf W}_{\text{time}}$ denote a $N_{\text{time}}\times(J+1)$ matrix of cubic spline functions ($J$ of which correspond to the cubic splines and 1 for an intercept), where $N_{\text{time}}=|\mathcal{T}|$, and let ${\bf W}_{\text{space}}$ denote a $N_{\text{space}}\times(1+M_{s}+M_{e}+M_{c})$ matrix of thin plate spline functions with an intercept, where $N_{\text{space}}=|\mathcal{S}|$. The design matrix $\bf W$} can then be written as
${\bf W}={\bf W}_{\text{space}}\otimes{\bf W}_{\text{time}}$ (E.5)
@@
(Here we suppose $N=N_{\text{space}}N_{\text{time}}$, although any inputs $\bf x$} with missing daily maximum temperature measurements are omitted.) The matrix $\bf W$} in Equation E.5 is then $N\times[(J+1)\times(1+M_{s}+M_{e}+M_{c})]$, meaning that this yields $(J+1)(1+M_{s}+M_{e}+M_{c})$ hyperparameters for the prior mean. We re-label the coefficients $\{\delta^{(\cdot)}_{jm}\}$ from Equation E.4 into an appropriately ordered vector $\boldsymbol{\beta}$} such that the prior mean function across all $N$ measurements can be concisely written as ${\bf W}{\boldsymbol{\beta}}$, as used in Section 3.1.
@@
Figure E.1: Comparison of the fitted prior mean (using ordinary least squares) versus input data for four selected stations (shown in panel a.) across all time points (panel b.), where the raw data are shown in black and the fitted prior mean in red. Panel c. then shows spatial maps of the 2-, 5-, and 10-day lagged autocorrelation from the input data (top) and residuals (data minus OLS prior mean; bottom), while panel d. shows the length-scale and signal variance from semivariograms fitted separately for each for the input data (top) and residuals (data minus OLS prior mean; bottom).
@@
In order to maintain a compromise between sufficient flexibility and overfitting, we fix $J=8$, $M_{s}=200$, $M_{e}=5$, and $M_{c}=5$, resulting in $1899$ hyperparameters for the prior mean. To visualize the selected prior mean function, we show the fitted prior mean – for starters using ordinary least squares (OLS), but later these hypereparameters will be trained online with the GP – for specific stations and days versus the corresponding raw temperature measurements; see Figure E.1. Across all input locations, the OLS linear mean explains 94.1% of the variability in the daily temperature data, which is impressive since there is one hyperparameter for every $\approx 1100$ data points. For the specific stations shown in Figure E.1(b), it is clear that the seasonal cycle appropriately captures the general trend of the data, such that the spatially-varying nature of the coeffients can describe different amplitudes and even shapes in the the seasonal cycle: note that the daily maximum temperatures range from approximately 5C to 35C at the Matador station and approximately -10C to 30C at the Sibley station; furthermore, the seasonal cycle is extremely smooth at the Minden station but “bumpy” at Morehead City.
@@
It is important to note how the variability explained by a flexible prior mean function “increases” the sparsity that can be leveraged by the kernel. In the time dimension, Figure E.1(c) shows how the 2-, 5-, and 10-day lagged autocorrelations at each gauged location differ for the input data versus the residuals (the input data minus the OLS prior mean function). Even at a lag of 10 days, the data show strong autocorrelations, in excess of 0.6. However, when removing the fitted OLS prior mean, 10-day lagged autocorrelations essentially vanish; even for 5-day lags, the autocorrelations are quite small. Spatially, Figure E.1(d) shows the length-scale and signal variance from a variogram fitted to either the input data (black) or the OLS residuals separately for each day. Here, the effect of the prior mean function is less clear: in particular, the length-scale of the residuals appear to be on a similar order of magnitude as the input data. On the other hand, the variability of the residuals is significantly reduced relative to the input data. In summary, it is clear that the prior mean function allows the data to be quite sparse in the time domain (i.e., correlations beyond $\pm 5$ days can be safely ignored) and perhaps less sparse in the spatial domain. These differences in the structure explained by the prior mean (temporal vs. spatial) have important implications for how we set up the kernel.
@@
Kernel. For the full space-time data, we use a core kernel similar to the one in Equation E.1, albeit generalized to a three-dimensional input space, i.e., $C_{\text{core}}({\bf x},{\bf x}^{\prime})$. The anisotropy matrix again uses the eigendecomposition ${\bf\Sigma}({\bf x})={\bf\Gamma}({\bf x}){\bf\Lambda}({\bf x}){\bf\Gamma}({\bf x})^{\top}$, where
${\bf\Lambda}({\bf x})=\left[\begin{array}[]{ccc}\lambda_{1}({\bf s})&0&0\\ 0&\lambda_{2}({\bf s})&0\\ 0&0&\lambda_{3}(t)\end{array}\right],\qquad{\bf\Gamma}({\bf x})=\left[\begin{array}[]{ccc}\cos\gamma({\bf s})&-\sin\gamma({\bf s})&0\\ \sin\gamma({\bf s})&\cos\gamma({\bf s})&0\\ 0&0&1\end{array}\right].$
@@
The formulas for $\lambda_{1}({\bf s})$, $\lambda_{2}({\bf s})$, and $\gamma({\bf s})$ are as in Equation E.2, and the length-scale in time is a constant, i.e., $\lambda_{3}(t)\equiv\lambda_{3}$. Note that in this case, the term inside the Matérn correlation is
@@
For the sparse kernel, the Wendland component is again anisotropic with coordinate-specific length scales (and coordinate-specific truncation limits); following Section 5.1 we use 10 units for the spatial truncation limits and 5 days for the temporal truncation limits based on the fact that the prior mean function explains autocorrelation beyond 5 days for a large majority of the domain. We again use 100 bump functions with $n_{1}=4$ and $n_{2}=25$; however, again based on the fact that the prior mean explains so much of the temporal variability, we restrict the bump functions to live in the spatial domain only. In total, this yields $521+1+1=523$ hyperparameters for the kernel (the same 521 from the spatial-only modeling in Section 5.1 plus $\lambda_{3}$ and the Wendland length-scale in the time dimension).
Appendix F Supplemental figures
Figure F.1: Sample bump functions $g({\bf x})$ from Equation 1 for different values of the shape parameter $b$ with fixed amplitude $a$ and radius $r$. As $b\rightarrow 0$, the bump function becomes uniform on $(-r,r)\subset\mathbb{R}^{1}$; as $b\rightarrow\infty$, the bump functions become delta function centered at $\mathbf{h}$. Note that the bump functions are exactly zero for distances from the centroid $\bf h$} greater than or equal to $r$.
Figure F.2: Root mean square error (RMSE), continuous rank probability score (CRPS), log-likelihood, and sparsity (defined as the fraction of the covariance that is nonzero) from analyzing the January 14, 2001 time slice of daily maximum temperature. We show results for the three configurations of $n_{1}$ and $n_{2}$ as well as the three upper bound limits on the Wendland length scale (“Wxx” indicates that the Wendland length scale is capped at $x$ units). The dashed gray line shows results from fitting a nonstationary Gaussian process with the same core kernel but without the sparsity-inducing kernel.
Figure F.3: Trace plots for the January 14, 2001 time slice of daily maximum temperature using the $n_{1}=4$, $n_{2}=25$ configuration.
Figure F.4: Trace plots for the January 14, 2001 time slice of daily maximum temperature using the $n_{1}=4$, $n_{2}=25$ configuration.
Figure F.5: Posterior summaries of hyperparameters for the January 14, 2001 time slice of daily maximum temperature. Panel (a) shows the posterior density of the prior mean; panels (b) and (c) show the square root of the signal variance and anisotropy length-scales, calculated using the posterior median hyperparameters and the physical covariates. Panel (d) shows implied correlations between eight reference locations and all other points in the input domain.
Figure F.6: Directional autocorrelation functions for pairs of locations over time. Here, we focus on four reference locations: Berkeley, CA; Minneapolis, MN; Nashville, TN; Portland, OR. For each of these reference locations, we compare correlation at four sites that are $2^{\circ}$ north, south, east, and west. In panel (a), the activated bump functions (i.e., those with $a_{ij}=1$) are shown in dashed gray. In panel (b), correlations are only plotted when they are nonzero.
Appendix G Supplemental tables
Parameter Description Default values
$n_{1}$ The number of bump function sums; defines the maximum possible groups of correlated, distance-unrelated regions. As large as possible, but realistically between 5 and 10. We found that $n_{1}\leq n_{2}$ resulted in better performance.
$n_{2}$ The number of bump functions in each $f_{i}$; defines the maximum possible regions of distance-unrelated correlations. As large as possible, but realistically between 10 and 50.
$D_{0}$ Prior upper bound for the compactly supported Wendland kernel. 5-15% of the maximum inter-point distance.
$D_{r}$ Prior upper bound for the bump function radii. 2-5% of the maximum inter-point distance.
$l_{k},u_{k}$ Prior lower and upper bound, respectively, for the bump function centroids in dimension $k=1,\dots,d$ The minimum and maximum values present in the data set.
Table G.1: Summary table of all user-specified parameters for the sparse kernel $C_{\text{sparse}}$, along with recommended default values.